OpenStudy (anonymous):

Solve: −5(x − 2) = 15

OpenStudy (anonymous):

First, you must distribute the -5 to x and -2: −5(x − 2) = 15 turns into -5x + 10 = 15 Now subtract 10: -5x + 10 = 15 turns into -5x=5 Now divide by -5: -5x=5 turns into x=5/-5, or x=-1 Do you understand this?